suppose, a solid sphere made of iron has a radius of . it travels 30cm in 3s in watar. density of iron= 7.5×10^3 kgm^-3 and density of water is 10^3 kgm^-3. coefficient of viscosity of water is 10^-3. wat is the force re I ingarding viscosity?

P.S. I am really sorry. My English is not good at all.
1
Expert's answer
2018-04-16T09:13:17-0400
The force of viscosity on a small sphere
F_d=6πηRv
The velocity
v=l/t
Thus
F_d=6×3.14×〖10〗^(-3)×0.01×0.3/3=1.88×〖10〗^(-5) N
Answer: 1.88×〖10〗^(-5) N